-2x-10y=-12
3x+y=4
find value of (x,y)

Answer by unlockmath(1688)

Hello,
Let's multiply the second equation by 10 so it will look like this:
-2x -10y= -12
(10)3x+(10)y=(10)4 Now we can add both equation (note: the y cancels out)
which leaves us with:
28x = 28
x = 1 Now plug 1 into either equation and you get
y = 1
Plug these into the equations and it works out.
RJ Toftness
www.math-unlock.com